Could you please help me solve this system?
10m-3n=5
5m+2n=-8
I thought I could solve it easily by the addition method, but I came up with some crazy fractions and got stuck. Please help.
This is how I started:
2(10m-3n=5)
3(5m+2n=-8)
20m-6n=10
15m+6n=-24
35m=-34
This is where I got stuck. Please help!
This problem did not come from a text book, it came from a worksheet.

Answer by Alan3354(69443)

10m-3n=5
5m+2n=-8
I thought I could solve it easily by the addition method, but I came up with some crazy fractions and got stuck. Please help.
This is how I started:
2(10m-3n=5)
3(5m+2n=-8)
20m-6n=10
15m+6n=-24
35m=-34 ****************
It's 35m = -14
m = -2/5
-----------
n = -3